OpenStudy (anonymous):

Line A intersects the y-axis at -6. Its slope is 7. What would be another point on the line?

OpenStudy (anonymous):

the equation of the line A is 7x-y=6. You can put a number instead of x or y and you will get the another coordinate.

OpenStudy (anonymous):

for example y=0 7x-0=6 than x=7/6 so the point is (7/6;0)

OpenStudy (anonymous):

A straight line has the equation \[y=mx+c\] where m is the slope (7 in your case) and the c is the intercept, i.e. where the line crosses the y-axis (-6 in your example). So the equation of the line in your case is \[y=7x-6\] To find out another point on the line, find out what the value of y is whenever x takes a certain value. For example, when \[x=2\], \[y=7*2-6=14-6=8\] So another point on this line is \[(x,y)=(2,8)\].